A hiker drops a stone from a 53.5 meter tall cliff. How long does it take the stone to reach the base of the cliff?
Answer by fractalier(6550)

The free fall equation is
y = (1/2)gt^2
so that
t = sqrt(2y / g)
t = sqrt(2(53.5)/9.8)
t = sqrt(10.92)
t = 3.3 seconds